# Strong anomalous diffusion for free-ranging birds

### v1

@@ V Discussion
Although SAND and bi-linear scaling have been previously reported, primarily in mathematical models, this is the first time SAND has been observed in ecological studies and specifically in animal movement. Importantly, our two stochastic models also reproduce SAND, although the different crossover points in adults and fledglings are reproduced only by the multi-mode model, suggesting that this model can be used to account for the observation. In particular, based on the parameterization of data to the multi-mode model, we found a higher degree of separation between modes in fledglings, compared to adults, such that a lower crossover point indicates greater inter-mode separation. Notably, our models do not reproduce the fact that fledglings do not display ballistic scaling. Our second main result is the unexpected transition from convex to concave behavior of the moment spectrum function, pointing to a typical time-scale of $\sim 5$ minutes and a closely related spatial home-range scale. Importantly, this transition was found for both adults and fledglings and was not previously observed in mathematical models. Here as well, our qualitative and quantitative results can be replicated using one of two stochastic models, both capturing key features of the empirical data.
To understand the significance of the 5-minute timescale, we note that the commuting velocity for these birds is $\sim 9$ m/s [55] and the maximum displacement per night is $\sim 2.5$ km as detailed above. Factoring both the maximum displacement and the velocity, a bird flying at $9$ m/s will cover 2.5 km at approximately 5 minutes. Therefore, based on this observation and the models detailed above we offer two interpretation of this timescale. The first is a coupling between the individual and the environment as supported by the BLW model: due to the finite home range of the bird, straight long-range flights are limited to $\sim 5$ minutes, above which interactions with the edges of the home range become important and inhibit the increase of the displacement. This explanation is supported by the similar radii of $R=2.5$ km in the BLW above, as well as by direct analysis of the home range of the birds, where in Appendix C and Fig. S5 [65] we show that indeed for the adult Barn Owls, the home range during breeding is $2-3$ km wide. On the other hand, we have found that – on average – the fledglings display significantly larger home ranges than the adults (Fig. S5 [65]). Another closely related explanation for the 5-minute timescale is based on the multi-mode correlated-velocity model, and does not require incorporating explicit boundaries as these emerge from the model. In the multi-mode model, mechanical or preferential behavior of the bird limits its commuting behavior, resulting in SAND at short times and successfully producing a transition from convex to concave $\lambda(q)$ without an explicit home range. Notably, the multi-mode model can result in a relatively short (average) daily maximum displacement but would allow for larger excursions, as found for the fledglings. Here, an important finding is that the straightforward parameterization of the model – without detailed fine-tuning – accurately reflects our qualitative results as well as the quantitative differences between adults and fledglings, suggesting that our two-state multi-mode model captures the dynamics to a significant degree.
Further relating the two stochastic models to the data, we suggest the following explanation for the emergence of SAND: in a heterogeneous landscape, an animal may alternate between an extensive commuting mode – search for resource-rich patches – and an intensive, area-concentrated searching mode – search for prey within a local patch [54]. Indeed, a major effort in the literature is to distinguish between local area-restricted searches and commutes [73, 74]. Thus, an analysis of unsegmented trajectories results in a mixture of directed flights and long stationary periods, which might resemble BLW dynamics [8, 54]. As discussed above, this mixture occurs at short times scales, up to $\sim 5$ minutes, either due to interactions with the environment – at long time-scales commutes are limited by the animal’s home-range – or due to properties of the behavioral states such as energy-consumption or food-availability. The BLW model thus acts as a simple model for SAND, simplifying the more ecologically inclined multi-mode dynamics.
@@
Finally, our findings highlight the significance of computing a continuous range of displacement moments when modeling animal movement. While MSD effectively captures behavior of mono-scaling processes, for behavior in general processes it may only capture the bulk of the displacement distribution, its tail, or can reside at the crossover between the two regimes. Thus, by integrating the entire spectrum of moments, we can more accurately test hypotheses in alignment with movement data, and develop stochastic simulations to predict movement patterns and gain deeper ecological insights. Moreover, the analysis of higher moments revealed distinctions between adults and fledglings directly relatable to their multi-mode and home-range behavior. As the key features observed above can be effectively modeled using two different models, an important finding of this manuscript is the connection between empirical data, the BLW model, and an ecologically oriented multi-mode model. The fact that both these general models reproduced empirical SAND suggests the prevalence of SAND across animal movement behaviors, extending beyond the movement of Barn Owls discussed here, and offering a new perspective on understanding ecological dynamics with multi-mode behavioral strategies.
VI Acknowledgments
@@
. For fieldwork and technical assistance we thank Y. Bartan, A. Levi, S. Margalit, R. Shaish, G. Rozman and other members of the Movement Ecology Lab and the Minerva Center for Movement Ecology. ATLAS development, maintenance, and studies have been supported by the Minerva Center for Movement Ecology, the Minerva Foundation, and ISF grants 965/15 and 1919/19 to R.N and S.T.; research on black-winged kites was supported also by JNF/KKL grant 60-01-221-18. R.N. also acknowledges support from Adelina and Massimo Della Pergola Chair of Life Sciences. E.B. acknowledges support of Israel Science Foundation’s grant 1614/21.
VII SUPPORTING INFORMATION
@@ Appendix C Area estimation
Supplementary Information
In this supplementary information we provide Figs. S1-S5 to support the discussion in the main text. In what follows, the notations and abbreviations are the same as in the main text and the equations and figures refer to those therein.
Figure S1: Results for the adult Barn Owls based on tracks with different durations: (a-c) 1-hour, (d-f) 2-hour and (g-i) 8-hour. (a,d,g) Log-log plot of the average absolute displacement moments $\langle|x(t)-x(0)|^{q}\rangle$. Data points for moment orders q=1 (blue), q=2 (orange), and q=3 (green) are shown. Circles represent empirical moments, while dashed lines indicate fits at short and long times. (b,e,h) $\lambda_{s}(q)$ as a function of $q$. The red triangles represent the empirical values over short time intervals and the dashed blue and orange lines represent the bi-linear fit slopes for low and high moments, respectively. Inset: $\lambda(q)$ for segmented commuting flights. (c,f,i) $\lambda_{\ell}(q)$ as a function of $q$ for long time intervals ($>5$ minutes, red triangles). The dashed blue and orange lines represent the bi-linear fit slopes for low and high moments, respectively.
Figure S2: Results for the fledgling Barn Owls. (a) Log-log plot of the average absolute displacement moments $\langle|\bm{x}|^{q}\rangle/\mathcal{R}_{0}^{q}$ with $\mathcal{R}_{0}=2.5$ km representing the average maximum displacement of Barn Owls in a single night. Data points for moment orders $q=1,1.5,2,2.5,3$ are shown, displaying two distinct scaling regimes for each moment $q$, with a transition at $\sim 5$ minutes. (b) $\lambda_{s}(q)$ as a function of $q$. The red triangles represent the empirical values derived from tracking data over short time intervals. The dashed blue and orange lines represent the bi-linear fit slopes for low and high moments, respectively, where the slopes are denoted in the legends. Inset: $\lambda(q)$ for segmented commuting flights, see text. (c) $\lambda_{\ell}(q)$ as a function of $q$ for long time intervals ($>5$ minutes, red triangles). The dashed blue and orange lines represent the bi-linear fit slopes for low and high moments, respectively.
Figure S3: Probability density function of jump times for adult and fledgling Barn Owls based on the segmentation detailed in the main text. The resulting distribution for both datasets are fitted to a power law and we find $\alpha\simeq 1.6$. The results are compared to BLW simulations ($x$ marks) with $\alpha=1.6$.
Figure S4: Results for the multi-mode simulations for the fledgling’s parameterization. (a) Log-log plot of the average absolute displacement moments $\langle|\bm{x}|^{q}\rangle/\mathcal{R}_{0}^{q}$ with $\mathcal{R}_{0}=2.5$ km representing the average maximum displacement of Barn Owls in a single night. Data points for moment orders $q=1,1.5,2,2.5,3$ are shown, displaying two distinct scaling regimes for each moment $q$, with a transition at $\sim 5$ minutes. (b) $\lambda_{s}(q)$ as a function of $q$. The red triangles represent the values over short time intervals of the simulations. The dashed blue and orange lines represent the bi-linear fit slopes for low and high moments, respectively, where the slopes are denoted in the legends. (c) $\lambda_{\ell}(q)$ as a function of $q$ for long time intervals ($>5$ minutes, red triangles). The dashed blue and orange lines represent the bi-linear fit slopes for low and high moments, respectively.
Figure S5: Area estimation: comparison between fledgling and adult Barn Owls. Panel (a) shows the home range of a fledgling Barn Owl, tagged as 4722, with its movements overlaying a satellite map. Panel (b) presents the home range for an adult Barn Owl, tagged as 4915. In both panels, the black dots represent the movement track, the red line is an Minimum Convex Polygon (MCP) estimation of the home range, and the yellow and purple lines indicate the 68th and 95th percentiles of the home range estimated using Kernel Density Estimation (KDE). Panel (c) compares the squared root of the home range areas – as inferred from the MCP – between fledglings and adults.

### v2

@@ VI Discussion
Another main result is the unexpected transition from convex to concave behavior of the moment spectrum function, pointing to a typical time-scale of $\sim 5$ minutes and a closely related spatial home-range scale. Importantly, this transition was found for both adults and fledglings and was not previously observed in mathematical models. Here as well, our qualitative and quantitative results can be replicated using one of two stochastic models, both capturing key features of the empirical data.
This is the first time bi-linear scaling and a transition from convex to concave scaling, have been observed in an ecological system, where most previous works have found SAND only in mathematical models. We therefore developed two stochastic models that largely reproduce these features, although the different crossover points ($q_{s}$ and $q_{\ell}$) in adults and fledglings are reproduced only by the multi-mode model, suggesting that this model can be used to better account for the observation. In addition, based on the parameterization of data to the multi-mode model, we found a higher degree of separation between modes in fledglings, compared to adults, such that a lower crossover point indicates greater inter-mode separation. Notably, our models do not fully reproduce the fledglings’ ballistic scaling (exponent $s_{2}=0.79$), and the near-saturation of moment spectrum function at long times (exponent $\ell_{2}=0.07$ for the adults). Nevertheless, we stress that the parameters of both the BLW and OU models were chosen independently – based on short-time scaling (BLW) or behavioral segmentation (OU) – and not optimized to fit the full empirical profile. The fact that these models reproduce key features of the data, including the crossover behavior and partial agreement with long-time scaling, underscores the interpretive value of our framework.
Importantly, many valuable approaches have been developed to analyze movement data, including step-length and turning-angle statistics, first-passage times, fractal dimensions, Hidden Markov Models, and composite or multi-scale movement models [16, 24, 95, 96, 97, 98, 99, 100, 101, 102, 103]. While SAND highlights some limitations of MSD-based analyses, it does not replace these existing methods. Rather, it offers a generalization of the MSD framework that captures moment-dependent and scale-dependent behavior. In parallel, we do not aim to cover the full range of ecological models; instead, we focus on two representative simulation classes – BLW and multi-mode correlated velocity models – to illustrate how SAND can extract structural features across diverse regimes.
@@
Finally, our findings highlight the significance of computing a continuous range of displacement moments when modeling animal movement. While MSD effectively captures behavior of mono-scaling processes, for behavior modeled by general processes it may only capture the bulk of the displacement distribution, its tail, or can reside at the crossover between the two regimes. Thus, by integrating the entire spectrum of moments, we can more accurately test hypotheses in alignment with movement data, and develop stochastic simulations to predict movement patterns and gain deeper ecological insights. Moreover, the analysis of higher moments revealed distinctions between adults and fledglings directly relatable to their multi-mode and home-range behavior. As the key features observed above can be effectively modeled using two different models, an important finding of this manuscript is the connection between empirical data, the BLW model, and an ecologically oriented multi-mode model. The fact that both these general models reproduced empirical SAND suggests the prevalence of SAND across animal movement behaviors, extending beyond the movement of Barn Owls discussed here, and offering a new perspective on understanding ecological dynamics with multi-mode behavioral strategies.
VII Acknowledgments
@@
For fieldwork and technical assistance we thank Y. Bartan, A. Levi, S. Margalit, R. Shaish, G. Rozman and other members of the Movement Ecology Lab and the Minerva Center for Movement Ecology. ATLAS development, maintenance, and studies have been supported by the Minerva Center for Movement Ecology, the Minerva Foundation, and ISF grants 965/15 and 1919/19 to R.N and S.T.. R.N. also acknowledges support from Adelina and Massimo Della Pergola Chair of Life Sciences. E.B. acknowledges support of Israel Science Foundation’s grant 1614/21.
VIII SUPPORTING INFORMATION